# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ 3 Propagation of chaos
For each $N\in\mathbb{N}$, let $(W^{i},Z^{i})$, $i=1,\ldots,N$, be independent copies of the pair $(W,Z)$. Denote by $N^{i}(dt,dz)$ the Poisson random measure corresponding to the jumps of the Lévy process $Z^{i}$ with intensity measure $\nu(dz)dt$, and define the compensated Poisson random measure by
@@
We now consider the system of non-interacting particles associated with the Lévy-driven McKean-Vlasov SDE with common noise (1.1), where the state $X^{i}=(X_{t}^{i})_{t\geq 0}$ of the $i$-th particle is defined by
@@
For $\bm{x}^{N}:=(x_{1},x_{2},\ldots,x_{N}),\bm{y}^{N}:=(y_{1},y_{2},\ldots,y_{N})\in\mathbb{R}^{dN}$, we have
$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\delta_{0})=\frac{1}{N}\sum_{i=1}^{N}|x_{i}|^{2}.$
@@
$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\mu^{\bm{y}^{N}})\leq\frac{1}{N}\sum_{i=1}^{N}|x_{i}-y_{i}|^{2}=\frac{1}{N}\left|\bm{x}^{N}-\bm{y}^{N}\right|^{2},$
(see inequality (1.24) in [18]). The true measure $\mathcal{L}_{X_{t}}$ at each time $t$ is approximated by the empirical measure
$\mu_{t}^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}({\rm{d}}x),$
@@
(see [41]) For each $p\in[2,p_{0}]$, there exists a constant $C_{p}>0$ such that, for all $t\geq 0$,
$\max_{1\leq i\leq N}\mathbb{E}\!\left[|X_{t}^{i,N}|^{p}\right]\leq\begin{cases}C_{p}\big{(}1+e^{\gamma pt}\big{)},&\gamma\neq 0,\\[3.00003pt] C_{p}(1+t)^{p/2},&\gamma=0,\ p=2\ \text{or}\ (\gamma=0,\ \gamma_{2}>0,\ p\in(2,p_{0}]),\\[3.00003pt] C_{p}(1+t)^{p},&\gamma=0,\ \gamma_{2}=0,\ p\in(2,p_{0}],\end{cases}$
@@
The argument follows the same reasoning as in the proof of Proposition 2.3 and is therefore not repeated here. ∎
Next, we state a result concerning the propagation of chaos, which plays a central role in the convergence analysis as $N\to\infty$. For clarity, we first introduce the following definition,
$\varphi(N)=\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\ln N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}$ (3.3)
(see Theorem 1 of [21]).
@@
Assume that all conditions in Proposition 3.1 hold and that Assumption (A2) holds for $L_{1}\in\mathbb{R}$, $L_{2}\geq 0$. Then, we have
$\max_{i\in\{1,\ldots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C_{T}\varphi(N),$ (3.4)
for any $N\in\mathbb{N}$, where the positive constant $C_{T}$ does not depend on $N$.
Assume further that $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$. Then, we have
$\max_{i\in\{1,\ldots,N\}}\sup_{t\geq 0}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C\varphi(N),$ (3.5)
for any $N\in\mathbb{N}$, where the positive constant $C$ does not depend on $N$ and $T$.
@@
$\quad+\int_{0}^{t}\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
@@
$e^{-\lambda t}|X_{t}^{i}-X_{t}^{i,N}|^{2}=\int_{0}^{t}e^{-\lambda s}\Big{[}-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\rangle$
@@
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}))z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
@@
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\left\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{X^{N}})\right\rangle\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}\left(|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}|c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm{d}}z){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{1}|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{X^{N}})\right){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
By taking the expectation and utilizing the bound
$\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}^{i}},\mathcal{L}^{1}_{X_{s}^{i,N}}\big{)}\right]\leq\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right],$
it follows that, for any $\varepsilon>0$,
@@
Moreover, by Proposition 3.1, for any $p\in(4,p_{0}]$, we have
@@
for some constant $C_{T}>0$. Following Fournier and Guillin [21] and Sznitman [38], for any $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ with $q>4$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that the empirical measure $\mu^{N}$ satisfies
$\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mu^{N},\mu)\right]\leq C\varphi(N),$
where $\varphi(N)$ is defined in (3.3). Although the original result in Fournier and Guillin [21] assumes independent particles, this bound still holds for our system with common noise $W^{0}$ due to the exchangeability of the particles conditioned on $\mathcal{F}^{0}$. The rest of the proof proceeds similarly by controlling the term $\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X^{i,N}_{s}},\mu^{X^{N}}_{s}\big{)}\right]$ using the triangle inequality and the exchangeability of the particles. If $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ for some $q>4$, that is, if
$M_{q}(\mu)^{q}:=\int_{\mathbb{R}^{d}}|x|^{q}\,\mu({\rm{d}}x)<\infty,$
then for each dimension $d\geq 1$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that, for all $N\geq 2$,
$\begin{split}\mathbb{E}\left[\mathcal{W}_{2}(\bar{\mu}^{N},\mu)^{2}\right]&\leq C\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\log N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}\\ &=C\varphi(N).\end{split}$
Next, let $(X_{t}^{i,N})_{i=1}^{N}$ be a collection of random variables that are identically distributed but not independent, and define the empirical measure by
$\mu_{t}^{\bm{x}^{N}}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i,N}}.$
Meanwhile, let $(X_{t}^{i})_{i=1}^{N}$ be a sequence of independent and identically distributed random variables, and define the corresponding empirical measure by
$\bar{\mu}_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i}}.$
In this case, we have
[…]
@@
$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\left(L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){{\rm{d}}}s\\ &\leq L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\left(6\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]+3C\varphi(N)\right){{{\rm{d}}}}s\\ &=6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.\end{split}$
Then using Gronwall’s inequality, we have
@@
$\begin{split}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]&\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)(1-e^{\lambda t})/\lambda}\\ &\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)/\lambda}\\ &=C\varphi(N),\end{split}$
for any $t\in[0,T]$, where the positive constant $C$ does not depend on time. Finally, when $\gamma<0$, it follows from Proposition 3.1 that for any $p\in(4,p_{0}]$,
$\max_{i\in\{1,\dots,N\}}\sup_{t\geq 0}\mathbb{E}\left[\left|X^{i,N}_{t}\right|^{q}\right]\leq C,$

### v2

@@ 3 Propagation of chaos
For each $N\in\mathbb{N}$, let $(W^{i},Z^{i})$, $i=1,\ldots,N$, be independent copies of the pair $(W,Z)$. Denote by $N^{i}({\rm{d}}t,{\rm{d}}z)$ the Poisson random measure corresponding to the jumps of the Lévy process $Z^{i}$ with intensity measure $\nu({\rm{d}}z){\rm{d}}t$, and define the compensated Poisson random measure by
@@
We now consider the system of non-interacting particles associated with the McKean-Vlasov SDE driven by Lévy processes with common noise (1.1), where the state $X^{i}=(X_{t}^{i})_{t\geq 0}$ of the $i$-th particle is defined by
@@
For $\bm{X}^{N}:=(X^{1},X^{2},\ldots,X^{N}),\bm{Y}^{N}:=(Y^{1},Y^{2},\ldots,Y^{N})\in\mathbb{R}^{dN}$, we have
$\mathcal{W}_{2}^{2}(\mu^{\bm{X}^{N}},\delta_{0})=\frac{1}{N}\sum_{i=1}^{N}|X^{i}|^{2}.$
@@
$\mathcal{W}_{2}^{2}(\mu^{\bm{X}^{N}},\mu^{\bm{Y}^{N}})\leq\frac{1}{N}\sum_{i=1}^{N}|X^{i}-Y^{i}|^{2}=\frac{1}{N}\left|\bm{X}^{N}-\bm{Y}^{N}\right|^{2},$ (3.3)
(see inequality (1.24) in [12]). The true measure $\mathcal{L}^{1}_{X_{t}}$ at each time $t$ is approximated by the empirical measure
$\mu_{t}^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}({\rm{d}}x),$ (3.4)
@@
([32, Proposition 3.1]) Let assumptions (A1)-(A7) hold, for each $p\in[2,p_{0}]$, there exists a constant $C_{p}>0$ such that, for all $t\geq 0$,
$\max_{1\leq i\leq N}\mathbb{E}\!\left[|X_{t}^{i,N}|^{p}\right]\leq\begin{cases}C_{p}\big(1+e^{\gamma pt}\big)&\gamma\neq 0,\\[3.00003pt] C_{p}(1+t)^{p/2}&\gamma=0,\ p=2\ \text{or}\ (\gamma=0,\ \gamma_{2}>0,\ p\in(2,p_{0}]),\\[3.00003pt] C_{p}(1+t)^{p}&\gamma=0,\ \gamma_{2}=0,\ p\in(2,p_{0}],\end{cases}$
@@
The argument follows the same reasoning as in the proof of Proposition 2.5 and is therefore omitted here. ∎
@@
Define the rate function
$\varphi(N)=\begin{cases}N^{-1/2}&\text{for }d<4,\\ N^{-1/2}\ln N&\text{for }d=4,\\ N^{-2/d}&\text{for }d>4.\end{cases}$ (3.6)
Under the conditions of Proposition 3.1 and Assumption (A2) with $L_{1}\in\mathbb{R}$, $L_{2}\geq 0$, there exists a constant $C_{T}>0$, independent of $N$, such that for all $N\in\mathbb{N}$,
$\max_{i\in\{1,\ldots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C_{T}\varphi(N).$ (3.7)
Furthermore, if $L_{1}+L_{2}<0$ and $\gamma<0$, then the constant can be made independent of time, i.e., there exists $C>0$, independent of $N$ and $T$, such that
$\max_{i\in\{1,\ldots,N\}}\sup_{t\geq 0}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C\varphi(N).$ (3.8)
@@
$+\int_{0}^{t}\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
@@
$e^{-\lambda t}|X_{t}^{i}-X_{t}^{i,N}|^{2}=\int_{0}^{t}e^{-\lambda s}\Big[-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\rangle$ (3.9)
@@
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\big(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\big)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
@@
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\left\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right\rangle\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}\left(|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}|c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm{d}}z){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg[\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg]$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{1}|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg[\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg]$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
By taking the expectation to (3.9), we have for any $\varepsilon>0$,
@@
Next, according to (3.2), (3.3), (3.4) and $\mathbb{E}\left[\mathcal{W}_{2}^{2}\big(\mathcal{L}^{1}_{X_{s}^{i}},\mathcal{L}^{1}_{X_{s}^{i,N}}\big)\right]\leq\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]$, let $|X_{s}^{j}-X_{s}^{j,N}|^{2}:=\max_{i\in\{1,\ldots,N\}}|X_{s}^{i}-X_{s}^{i,N}|^{2}$ we have
$\begin{split}\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]&\leq\mathbb{E}[\mathcal{W}_{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mathcal{L}^{1}_{X_{s}^{i}})+\mathcal{W}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\mathcal{W}_{2}(\mu^{\bm{X}^{N}},\mu_{s}^{\bm{X}^{N}})]^{2}\\ &\leq 3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mathcal{L}^{1}_{X_{s}^{i}})+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\mathcal{W}^{2}_{2}(\mu^{\bm{X}^{N}},\mu_{s}^{\bm{X}^{N}})]\\ &\leq 3\mathbb{E}[\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i}-X_{s}^{i,N}|^{2}]\\ &\leq 3\mathbb{E}[\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+|X_{s}^{j}-X_{s}^{j,N}|^{2}]\\ &\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})].\end{split}$ (3.11)
Moreover, by Proposition 3.1, for any $p\in[2,p_{0}]$, we have
@@
for some constant $C_{T}>0$, together with Carmona and Delarue [10, Theorem 5.8] deduce that
$\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})]\leq C\varphi(N).$ (3.12)
Combining (3.11) and (3.12), we have
$\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]$ $\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})]$ (3.13)
$\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3C\varphi(N).$
@@
$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\left(L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){{\rm{d}}}s\\ &\leq L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\left(6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3C\varphi(N)\right){{{\rm{d}}}}s\\ &=6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.\end{split}$ (3.14)
Thus, we have
$e^{-\lambda t}\mathbb{E}\left[|X_{t}^{j}-X_{t}^{j,N}|^{2}\right]\leq 6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.$
Then using Grönwall’s inequality, we have
@@
$\begin{split}\mathbb{E}\left[|X_{t}^{j}-X_{t}^{j,N}|^{2}\right]&\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)(1-e^{\lambda t})/\lambda}\\ &\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)/\lambda}\\ &=C\varphi(N),\end{split}$